# Beyond Homophily: Community Search on Heterophilic Graphs

### v1

@@ VIII-A Experimental Setup
(a) Efficiency results of the training phase (per epoch)
@@
VIII-B Effectiveness of AdaptCS
Table II benchmarks our approach on 16 graphs, where AdaptCS-I is the vanilla model that performs distance- and frequency-aware aggregation on the full adjacency; AdaptCS-II adds the proposed low-rank SVD compression, which keeps hop-distinct features in the latent space and thus avoids memory overflow. For the online search phase, we compare two method: the Signed Community Search (SCS) and Adaptive Community Score (ACS). “-” indicates that the method exceeds the GPU memory limit or fails to finish within the 12-hour wall-clock time. The rightmost column, Ave +/-, reports the average F1-score difference of each method relative to the best AdaptCS-II (ACS) result on average across all datasets.
Observation 3 – Homophily-oriented GNNs underperformed. Among the four GNN backbones, QDGNN attains the best overall averages, followed by ComDF, then ICSGNN, while COCLEP trails behind. Across the three heterophily extensions, Tanh delivers the largest average gains, ACM offers the second-best but slightly smaller improvements, and ALT is the least effective with the highest computation overhead. Even under the strongest extension, the averages still fall short of AdaptCS, indicating that flip effects and distance-agnostic aggregation remain unresolved in these backbones.
VIII-C Efficiency of AdaptCS
Figure 4 reports the training time of all learning-based models and the query efficiency of all search algorithms. ACS shares the same offline time as SCS, so it is omitted for clarity.
Training phase. SVD compression eliminates OOM and shortens runtime. AdaptCS-II is the only model that fits into a single GPU on the 110-million-edge Reddit graph, finishing the offline stage within 40s, whereas every other GNN variant crashes with OOM. AdaptCS-II achieves similar or better runtime compared to AdaptCS-I. Compared to baselines, both AdaptCS variants demonstrate better training efficiency, especially for the ACM-enhanced baselines, which pay an additional cost for channel mixing. The numbers in Figure 4 report per-epoch training time and therefore exclude the one-off SVD pre-computation; the pre-computation takes 0.2s on Texas and up to 90s on Reddit. This cost can be ignored after being amortized over hundreds of training epochs.
Query phase. AdaptCS-SCS achieves sub-millisecond latency on most graphs, while AdaptCS-ACS achieves the best efficiency and stays below 0.03s on large graphs. Compared with QDGNN or COCLEP, which require several seconds of forward propagation per query on heterophilic graphs such as Film, both AdaptCS objectives achieve up to 2 orders of magnitude acceleration. Classical heuristics (clique, $k$-truss, $k$-core) return results in microseconds on small graphs; ACS matches the efficiency of classical heuristics on large graphs.
To sum up, the SVD-optimized AdaptCS-II not only attains the best accuracy but also solves the memory bottleneck: it is the only method that completes training on all datasets, including Flickr and Reddit, and its query latency remains competitive with lightweight heuristics while outperforming all learning-based baselines by large margins.
@@ VIII-D Ablation Study
Adaptive masking. Replacing the adaptive masking with hard masking (w/o mask) results in the most significant performance drop across most datasets, confirming that isolating exact-$k$ neighborhoods while retaining triangle-rich edges is fundamental for model effectiveness.
@@
MLP vs. attention. Our framework supports both MLP-based fusion and per-class attention bank for hop channel mixing. In practice, both approaches achieve comparable performance across benchmarks, with MLP-based fusion showing slightly better results on larger datasets. This improvement is likely due to the greater expressive power of MLPs, which can capture complex relationships and adapt to the structural diversity present on large-scale graphs.
@@ VIII-E Hyper-parameter Analysis
Similarity threshold $\boldsymbol{\tau}$6(a). Similarity threshold serves the purpose of transforming the cosine similarity of the learned embeddings into a signed graph; an edge $(u,v)$ is labeled positive if $s_{uv}\geq\tau$ and negative if $s_{uv}<\tau$. We test the threshold range of $\tau\!\in\!\{0.2,0.5,0.6,0.8,0.9\}$. From the study, we find that scores remain stable between $0.5\!\leq\!\tau\!\leq\!0.8$ and peak at $\tau\!=\!0.9$. A very small threshold ($\tau\!=\!0.2$) falsely turns many weakly related pairs into positive edges, whereas a strict threshold ($\tau\!=\!0.9$) may disconnect the signed graph and assign more weights to nodes that demonstrate high query similarity globally. The method is tolerant of moderate changes, and we retain $\tau\!=\!0.9$ as the default setting for experiments.
SVD rank $\boldsymbol{r}$6(b). The SVD rank $r$ determines the number of leading spectral components used in the low-rank approximation. Across most datasets, F1-score remains largely stable as $r$ increases from 16 up to 128, indicating that the low-rank operator is robust and most structural information is preserved even at low rank. Notably, a clear improvement is observed at $r=100$, where performance peaks or reaches its best on several datasets. Based on this observation, we set $r=100$ as the default, with $r=128$ used only for the largest graph (Reddit) to ensure maximum coverage.
Hop number $\boldsymbol{k}$6(c). The hop count decides how many exact-hop channels are aggregated, balancing receptive fields against noise. We vary $k$ between 3 and 9. Performance improves until $k$ reaches 5, after which additional hops add marginal information and even lower scores on sparse graphs such as Chameleon. Because long-range signals introduce weakly related nodes, we fix k=5k\!=\!5 in all experiments.
@@
IX Conclusions
In this study, we introduce AdaptCS, a novel community search framework adaptively designed for both homophilic and heterophilic graphs. Distance awareness through distinctive-hop aggregation and frequency awareness via low- and high-pass filtering, the AdaptCS encoder produces embeddings that remain discriminative even when neighboring nodes are largely dissimilar. A scalable low-rank SVD optimization further removes the memory bottleneck of high-order adjacency computation, enabling efficient training on graphs with over one hundred million edges. Leveraging these embeddings, the Adaptive Community Score (ACS) balances embedding similarity and topological relations, supporting accurate and efficient query-time retrieval. Extensive experiments on both heterophilic and homophilic benchmarks demonstrate that AdaptCS consistently outperforms state-of-the-art baselines, maintains robustness under varying degrees of heterophily, and achieves substantial gains in computational efficiency.
X Statement on AI-assisted tools

### v2

@@ VIII-B Effectiveness of AdaptCS
(a) Efficiency results of the training phase
@@
VIII-C Efficiency of AdaptCS
Figure 4 reports the training time of all learning-based models and the query efficiency of all search algorithms. ACS shares the same offline time as SCS, so it is omitted for clarity. The one-time offline cost (e.g., model training or preprocessing) and the online per-query latency are reported separately: Fig. 4(a) measures the offline training time for learning-based models, while Fig. 4(b) measures the online query time per query after the offline stage is completed.
Training phase. For a fair comparison, we report the offline training time of all learning-based baselines under the same light heterophily extension (ACM). SVD compression eliminates OOM and shortens runtime. The numbers reported for AdaptCS training time included the one-off SVD pre-computation cost. Overall, AdaptCS-II is the only model that fits into a single GPU on the 110-million-edge Reddit graph, whereas every other GNN variant crashes with OOM. On small graphs, AdaptCS-II has comparable training time to AdaptCS-I, while on large graphs, it is faster due to the SVD approximation. Compared to baselines, both AdaptCS variants demonstrate better training efficiency.
Query phase. The considered training extensions (e.g., ALT/ACM/TANH) only affect the offline training stage and do not change the online search procedure; therefore, the per-query efficiency reported below is identical across these extensions for the same model. AdaptCS-SCS achieves sub-millisecond latency on most graphs, while AdaptCS-ACS achieves the best efficiency and stays below 0.03s on large graphs. Compared with QDGNN or COCLEP, which require several seconds of forward propagation per query on heterophilic graphs such as Film, both AdaptCS objectives achieve up to 2 orders of magnitude acceleration over the strongest ML-based CS baselines. Classical heuristics (clique, $k$-truss, $k$-core) return results in microseconds on small graphs, and ACS matches their efficiency on large graphs.
To sum up, the SVD-optimized AdaptCS-II not only attains the best accuracy but also solves the memory bottleneck: it is the only method that completes training on all datasets, and its query latency remains competitive with heuristics while outperforming all learning-based baselines by large margins.
@@ VIII-D Ablation Study
Adaptive masking. Replacing the adaptive masking with hard masking (w/o mask) results in the significant drop across most datasets, confirming that isolating exact-$k$ neighbors while retaining triangle-rich edges is crucial for model performance.
@@
MLP vs. attention. Our framework supports both MLP-based fusion and a per-class attention bank for hop channel mixing. Overall, both fusion strategies are effective, while the attention-based fusion is stronger on most benchmarks; the MLP variant remains competitive and achieves comparable results on several datasets. We attribute the advantage of attention to its node- and class-adaptive weighting of hop channels, which better aligns the fused representation with query-relevant semantics, especially under heterophily.
@@ VIII-E Hyper-parameter Analysis
Similarity threshold $\boldsymbol{\tau}$6(a). Similarity threshold serves the purpose of transforming the cosine similarity of the learned embeddings into a signed graph; an edge $(u,v)$ is labeled positive if $s_{uv}\geq\tau$ and negative if $s_{uv}<\tau$. We test the threshold range of $\tau\!\in\!\{0.2,0.5,0.6,0.8,0.9\}$. From the study, we find that scores remain stable between $0.5\!\leq\!\tau\!\leq\!0.8$ and peak at $\tau\!=\!0.9$. A very small threshold ($\tau\!=\!0.2$) falsely turns many weakly related pairs into positive edges, whereas a strict threshold ($\tau\!=\!0.9$) assign more weights to nodes that demonstrate high query similarity globally. We retain $\tau\!=\!0.9$ as the default setting for experiments. Since similarity score distributions can vary across graphs, $\tau$ can also be set in a graph-adaptive manner, e.g., by choosing $\tau$ as the $p$-quantile of sampled edge similarities. This provides a simple option to stabilize the positive-edge ratio on unseen graphs.
SVD rank $\boldsymbol{r}$6(b). The SVD rank $r$ determines the number of leading spectral components used in the low-rank approximation. Notably, a clear improvement is observed at $r=100$, where performance peaks or reaches its best on several datasets. Based on this observation, we set $r=100$ as the default, with $r=128$ used only for the largest graph (Reddit) to ensure maximum coverage.
Hop number $\boldsymbol{k}$6(c). The hop count decides how many exact-hop channels are aggregated, balancing receptive fields against noise. We vary $k$ between 3 and 9. Performance increases as $k$ grows up to $5$, while further hops provide diminishing gains and can hurt accuracy on sparse graphs such as Chameleon. Since long-range propagation tends to introduce weakly related nodes, we set $k=5$ for all experiments.
@@
IX Case Study
Figure 7: Case Study: Distribution of Triangle Support
Figure 7 compares the triangle support of direct edges that are retained in 2-hop as strengthened by the adaptive mask ($S^{+}$) against those that are not ($S^{-}$). Across datasets, the $S^{+}$ distribution shifts to the right, evidencing heavier tails and larger means (e.g., Cora: $68.7$ vs. $18.4$; Squirrel: $215.8$ vs. $82.8$), which indicates that adaptively retained edges participate in more triangles than their non-retained counterparts.
Implication. Edges with higher triangle support anchor dense subgraphs: multiple common neighbors provide more two-hop paths that (i) concentrate local cohesiveness within the same semantic group, making endpoints more likely to belong to the same community, and (ii) improve robustness under heterophily, since cross-class or noisy links are diluted by several consistent neighbors. Hence, adaptive masking preserves triangle-rich edges that benefit community identification.
X Conclusions
In this study, we introduce AdaptCS, a novel community search framework adaptively designed for both homophilic and heterophilic graphs. By incorporating distance awareness through distinctive-hop aggregation and frequency awareness via low- and high-pass filtering, the AdaptCS encoder produces embeddings that remain discriminative even when neighboring nodes are largely dissimilar. A scalable low-rank SVD optimization further removes the memory bottleneck of high-order adjacency computation, enabling efficient training on graphs with over one hundred million edges. Leveraging these embeddings, the Adaptive Community Score (ACS) balances embedding similarity and topological relations, supporting accurate and efficient query-time retrieval. Extensive experiments on both heterophilic and homophilic benchmarks demonstrate that AdaptCS consistently outperforms state-of-the-art baselines, maintains robustness under varying degrees of heterophily, and achieves substantial gains in efficiency.
XI Statement on AI-assisted tools